OpenStudy (anonymous):

Simplify, using the properties of rational exponents. y2/3 ------ y1/3

OpenStudy (muzzack):

could you do a screenshot

OpenStudy (anonymous):

nope im on computer

OpenStudy (anonymous):

\[\frac{ y ^{2/3} }{ y ^{1/3} }\]

OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

When you divide, you subtract the exponents so you have, \[y ^{2/3-1/3}\]

OpenStudy (anonymous):

y1

OpenStudy (anonymous):

2/3 - 1/3 =?

OpenStudy (anonymous):

y1/0

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }-\frac{ 1 }{ 3 } = \frac{ 1 }{ 3 } \implies y ^{1/3}\]

OpenStudy (anonymous):

Please go over your fractions it will be very helpful.